# Task-Agnostic Machine Learning-Assisted Inference

### v1

@@ A.2 Proof the Proposition 1
where $\bm{\psi}(\cdot,\cdot;\bm{\theta})$ is a user-defined function. By the theory of Z-estimation and recent paper on ML-assisted inference (Miao et al., 2023b, ), we have ${\bf V}(\widehat{\bm{\theta}}_{\cal L})={\bf A}_{\bm{\theta}^{*}}^{-1}{\bf M}_{1}{\bf A}_{\bm{\theta}^{*}}^{-1},{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})={\bf A}_{\bm{\eta}}^{-1}{\bf M}_{4}{\bf A}_{\bm{\theta}^{*}}^{-1},{\bf V}(\widehat{\bm{\eta}}_{\cal L})={\bf A}_{\bm{\eta}}^{-1}{\bf M}_{2}{\bf A}_{\bm{\eta}}^{-1}$, and ${\bf V}(\widehat{\bm{\eta}}_{\cal U})={\bf A}_{\bm{\eta}}^{-1}{\bf M}_{3}{\bf A}_{\bm{\eta}}^{-1}$. Here, ${\bf M}_{1}=\textnormal{Var}_{l}[\bm{\psi}(Y,{\bf X};\bm{\theta}^{*})],{\bf M}_{2}=\textnormal{Var}_{l}[\bm{\psi}(\widehat{f},{\bf X};\bm{\theta}^{*})],{\bf M}_{3}=\textnormal{Var}_{u}[\bm{\psi}(\widehat{f},{\bf X};\bm{\theta}^{*})],{\bf M}_{4}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};\bm{\theta}^{*}),\bm{\psi}(\widehat{f},{\bf X};\bm{\theta}^{*})],{\bf A}_{\bm{\theta}^{*}}=\mathbb{E}[\partial\bm{\psi}(Y,{\bf X};\bm{\theta}^{*})/\partial{\bm{\theta}}],{\bf A}_{\bm{\eta}}=\mathbb{E}[\partial\bm{\psi}(\widehat{f},{\bf X};\bm{\eta})/\partial\bm{\eta}]$, and $\rho=\dfrac{n}{N}$.
@@
$\bm{\omega}=\bm{\omega}_{0}$ $=(\operatorname{Var}(\widehat{\bm{\eta}}_{\cal L})+\frac{n}{N}\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U}))^{-1}\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ (53)
@@
By the theory of POP-Inf, PPI++, and PPI paper (Miao et al., 2023b, ; Angelopoulos et al., 2023b, ; Angelopoulos et al., 2023a, ), we have
$n^{1/2}(\widehat{\bm{\theta}}_{\texttt{POP-Inf}}-\bm{\theta}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},\bm{\Sigma}({\rm diag}(\bm{\omega}^{*}_{\text{ele}})))$ (63)
@@
Denote ${\bf C}={\bf A}_{\bm{\theta}^{*}}^{-1}{\bf A}_{\bm{\eta}}$, we have POP-Inf, PPI++, and PPI is asymptotically equivalent to ${\rm diag}(\bm{\omega}^{*}_{\text{ele}}){\bf C},{\rm diag}(\bm{\omega}^{*}_{\text{tr}}){\bf C}$, and ${\rm diag}({\bf 1}){\bf C}$, respectively. This completes the proof. ∎
@@ A.3 Proof of Proposition 2
$n^{1/2}\left(\begin{array}[]{c}\widehat{\bm{\theta}}_{{\cal L}}-\bm{\theta}\\ \widehat{\bm{\eta}}_{{\cal L}}-\bm{\eta}\\ \widehat{\bm{\eta}}_{{\cal U}}-\bm{\eta}\\ \end{array}\right)\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left\{\left(\begin{array}[]{c}{\bf 0}_{K}\\ {\bf 0}_{K}\\ {\bf 0}_{K}\end{array}\right),\left(\begin{array}[]{ccc}{\bf V}(\widehat{\bm{\theta}}_{\cal L})&{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf 0}\\ {\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf V}(\widehat{\bm{\eta}}_{\cal L})&{\bf 0}\\ {\bf 0}&{\bf 0}&\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U})\end{array}\right)\right\},$ (82)
@@
A.4 Proof of Proposition 3
We state the following definition and lemma that will be used later.
Definition 1 (positive regression dependence on the subset (PRDS)).
For any increasing set $D$, and for each $i\in I_{0},P(\mathbf{X}\in D\mid\left.X_{i}=x\right)$ is non-decreasing in $x$.
Lemma 1.
(Benjamini and Yekutieli,, 2001) If the joint distribution of the test statistics is PRDS on the subset of test statistics corresponding to true null hypotheses, the BH procedure controls the FDR at level less than or equal to $\dfrac{K_{0}}{K}q$.
This lemma is Theorem 1.2 in Benjamini and Yekutieli, (2001), thus the proof is omitted.
We first show that the asymptotic p-value for POP-LM can be derived from the asymptotic distribution of the test statistics. By theorem 1, we have
$n^{1/2}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}-\bm{\theta}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left(0,{\bf V}(\widehat{\bm{\theta}}_{\cal L})-{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})\right),$ (84)
The asymptotic valid p-value can then be calculated by the test statistics
$\dfrac{n^{1/2}\widehat{{\bm{\theta}}}_{\textnormal{{PSPS}}_{k}}}{\operatorname{Var}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})_{kk}}\sim\mathcal{N}\left(0,1\right),$ (85)
By the PRDS property on the p-value, applying lemma 1 completes the proof.
A.5 Proof of Proposition 4
This proof is based on (Barber and Candès,, 2019). We state the following definition and lemma that will be used later.
Lemma 2.
(Barber and Candès,, 2019) Suppose that $Z_{1},\ldots,Z_{K}$ are independent with $Z_{k}\sim\operatorname{Bernoulli}\left(\rho_{k}\right)$, and $\min_{k}\rho_{k}\geq\rho>0$. Let $J$ be a stopping time in reverse time with respect to the filtration $\left\{\mathcal{F}_{k}\right\}$, where
$\mathcal{F}_{k}=\sigma\left(\left\{\left(Z_{1}+\cdots+Z_{k}\right),Z_{k+1},\cdots,Z_{d}\right\}\right),$ (86)
with $\sigma(\cdot)$ denoting a $\sigma$-algebra. Then
$\mathbb{E}\left[\frac{1+K}{1+Z_{1}+\cdots+Z_{K}}\right]\leq\rho^{-1}.$ (87)
Define the $S_{k}=\textnormal{sign}(W_{k})\in\{-1,0,1\}$, a random subset of $\mathcal{N}$ as $\mathcal{M}:=\mathcal{N}\backslash\left\{k\in\mathcal{N}:C_{k}=0\right\}=\{k\in\mathcal{N}:S_{k}\neq 0$. Here, we consider the difference statistic
$W_{k}=\left|\hat{\beta}_{k}(\lambda)\right|-\left|\hat{\beta}_{k+K}(\lambda)\right|,k=1,\ldots,K,$ (88)
where $\hat{\beta}_{k}(\lambda)$ and $\hat{\beta}_{k+K}(\lambda)$ is the debiased Lasso coefficient for the true and knockoff variables, respectively.
The FDR is equal to
FDR $=\mathbb{E}\left[\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\geq\tau^{\textnormal{knockoff}}\right\}}{\#\left\{k:W_{k}\geq\tau^{\textnormal{knockoff}}\right\}\vee 1}\right]$ (89)
$=\mathbb{E}\left[\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\geq\tau^{\textnormal{knockoff}}\right\}}{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}\cdot\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}{\#\left\{k:W_{k}\geq\tau^{\textnormal{knockoff}}\right\}\vee 1}\right]$ (90)
$\leq\mathbb{E}\left[\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\geq\tau^{\textnormal{knockoff}}\right\}}{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}\cdot\dfrac{\#\left\{k:W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}{\#\left\{k:W_{k}\geq\tau^{\textnormal{knockoff}}\right\}\vee 1}\right]$ (91)
$\leq q\cdot\mathbb{E}\left[\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\geq\tau^{\textnormal{knockoff}}\right\}}{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}\right]$ (92)
We reordered the indices of the null $W_{k}^{\prime}s$ so that $\left|W_{(1)}\right|\geq\left|W_{(2)}\right|\geq\ldots$. Let $Z_{k}={\bf 1}_{W_{(k)}}<0$. Let $K$ be the index such that $\left|W_{(1)}\right|\geq\cdots\geq\left|W_{(K)}\right|\geq\tau^{\textnormal{knockoff}}>\left|W_{(K+1)}\right|\geq\cdots$.
For each $k$, we have
$W_{(k)}\geq\tau^{\textnormal{knockoff}}\quad$ $\Leftrightarrow\quad\left|W_{(k)}\right|\geq\tau^{\textnormal{knockoff}}\text{ and }W_{(k)}>0\quad\Leftrightarrow\quad j\leq J\text{ and }Z_{k}=0\text{,}$ (94)
$W_{(k)}\leq-\tau^{\textnormal{knockoff}}\quad$ $\Leftrightarrow\quad\left|W_{(k)}\right|\geq\tau^{\textnormal{knockoff}}\text{and }W_{(k)}<0\quad\Leftrightarrow\quad j\leq J\text{ and }Z_{k}=1$ (95)
Therefore, we have
$\mathbb{E}\left[\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\geq\tau^{\textnormal{knockoff}}\right\}}{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}\right]$ $=\frac{\left(1-Z_{1}\right)+\cdots+\left(1-Z_{k}\right)}{1+Z_{1}+\cdots+Z_{k}}$ (96)
$=\frac{1+K}{1+Z_{1}+\cdots+Z_{k}}-1.$ (97)
By using debiased Lasso coefficient as to construct the $W_{k}$, we have asymptotically $Z_{k}\stackrel{{\scriptstyle\text{ i.i.d. }}}{{\sim}}$ Bernoulli(0.5). With $\rho=0.5$ in lemma 2, we have
$\mathbb{E}\left[\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\geq\tau^{\textnormal{knockoff}}\right\}}{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}\right]\leq 1$ (98)
Therefore,
$\textnormal{FDR}=q\cdot\mathbb{E}\left[\dfrac{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\geq\tau^{\textnormal{knockoff}}\right\}}{\#\left\{k:\beta_{k}=0\text{ and }W_{k}\leq-\tau^{\textnormal{knockoff}}\right\}+1}\right]\leq q,$ (99)
which completes the proof.
Appendix B Simulation details
@@
Here, we provide the details for our simulation. All our simulation is run in R with version 4.2.1 (2022-06-23) in a macbook air with M1 chip. For all the simulation, the ground truth coefficients are obtained using $5\times 10^{4}$ samples; A pre-trained random forest with 500 trees to grow is obtained from a hold-out data. We bootstrap the labeled data for $200$ times for covariance estimation. All simulations are repeated 1000 times.
@@
  • Mean estimation, Linear regression, and Quantile regression: We simulate the data from $Y_{i}=X_{1i}\beta_{1}+X_{2i}\beta_{2}+X_{1i}^{2}\beta_{2}+X_{2i}^{2}\beta_{2}+\epsilon_{i}$, where $X_{i1}$ and $X_{2i}$ are independent simulated from $\mathcal{N}(0,1)$, $\beta_{1}=\sqrt{0.08}$, $\beta_{2}$ is set to be the value such that $X_{2i}\beta_{2}+X_{1i}^{2}\beta_{2}+X_{2i}^{2}\beta_{2}$ explains $81\%$ of the outcome variance, and $\epsilon_{i}$ is simulated from a mean zero normal distribution with variance such that $\operatorname{Var}(Y_{i})=1$. We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with samples size $1000,2500,5000$, or $10000$.
@@
  • Logistic regression: We simulate the data from $Y_{i}=\mathds{1}(\tilde{Y}_{i}>\text{ median }(\tilde{Y}_{i}))$, where $\tilde{Y}_{i}=X_{1i}\beta_{1}+X_{2i}\beta_{2}+X_{1i}^{2}\beta_{2}+X_{2i}^{2}\beta_{2}+\epsilon_{i}$, where $X_{i1}$ and $X_{2i}$ are independent simulated from $\mathcal{N}(0,1)$, $\beta_{1}=\sqrt{0.08}$, $\beta_{2}$ is set to be the value such that $X_{2i}\beta_{2}+X_{1i}^{2}\beta_{2}+X_{2i}^{2}\beta_{2}$ explains $81\%$ of the outcome variance, and $\epsilon_{i}$ is simulated from a mean zero normal distribution with variance such that $\operatorname{Var}(\tilde{Y}_{i})=1$. We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with samples size $1000,2500,5000$, or $10000$.
@@
    We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with samples size $1000,2500,5000$, or $10000$. The $Z_{i}$ is used as a instrument for $X_{1i}$.
@@
    We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with samples size $1000,2500,5000$, or $10000$.
@@
    We use $X_{1i},\dots,X_{200i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 100 individuals, and unlabeled data with samples size $1500,2000,2500$, or $3000$.
@@
    where $\beta_{1}=\sqrt{0.01}$ to power simulation and $\beta_{1}=0$ for type-I error simulation, $\beta_{2}$ is set to be the value such that $\beta_{2}X_{2i}+\beta_{2}X_{2i}^{2}$ explains $81\%$ of the outcome variance, and $\tau_{\epsilon}$ is set to be the value such that $\operatorname{Var}(Y_{i})=0$.We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with samples size $1000,2500,5000$, or $10000$.
@@
    The correlation $r$ is set to be 0.25. We then simulate the outcome $Y_{i}=\sum_{k=1}^{150}X_{ki}\beta_{k}+\epsilon_{i}$, where we randomly sample $15$ $\beta_{k}$ to be 0.15 and let all other remaining $\beta_{k}$ to be 0. $\epsilon_{i}$ is simulated from a mean-zero normal distribution with variance set to the value such that $\operatorname{Var}(Y_{i})=1$. We further generate $Z_{i}=0.9Y_{i}+\sum_{k=1}^{150}X_{ki}\theta_{k}+\gamma_{i}$, where $\theta_{k}=0.15$ for all $k=1,\dots,150$. We use $Z_{i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with samples size $5000$.
  • knockoff: We used the same setting as described in the BH procedure above to generate the data. The only difference is that we set $beta_{k}=0.5$ for features associated with the outcome and considered labeled data consisting of 100 individuals, along with unlabeled data comprising a sample size of 1000.

### v3

@@ A.2 Proof the proposition 1
where $\bm{\psi}(\cdot,\cdot;\bm{\theta})$ is a user-defined function. By the theory of Z-estimation and a recent paper on ML-assisted inference [35], we have ${\bf V}(\widehat{\bm{\theta}}_{\cal L})={\bf A}_{{\bm{\theta}}^{*}}^{-1}{\bf M}_{1}{\bf A}_{{\bm{\theta}}^{*}}^{-1},{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})={\bf A}_{\bm{\eta}}^{-1}{\bf M}_{4}{\bf A}_{{\bm{\theta}}^{*}}^{-1},{\bf V}(\widehat{\bm{\eta}}_{\cal L})={\bf A}_{\bm{\eta}}^{-1}{\bf M}_{2}{\bf A}_{\bm{\eta}}^{-1}$, and ${\bf V}(\widehat{\bm{\eta}}_{\cal U})={\bf A}_{\bm{\eta}}^{-1}{\bf M}_{3}{\bf A}_{\bm{\eta}}^{-1}$. Here, ${\bf M}_{1}=\textnormal{Var}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}^{*})],{\bf M}_{2}=\textnormal{Var}_{l}[\bm{\psi}(\widehat{f},{\bf X};{\bm{\theta}}^{*})],{\bf M}_{3}=\textnormal{Var}_{u}[\bm{\psi}(\widehat{f},{\bf X};{\bm{\theta}}^{*})],{\bf M}_{4}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}^{*}),\bm{\psi}(\widehat{f},{\bf X};{\bm{\theta}}^{*})],{\bf A}_{{\bm{\theta}}^{*}}=\mathbb{E}[\partial\bm{\psi}(Y,{\bf X};{\bm{\theta}}^{*})/\partial{\bm{\theta}}],{\bf A}_{\bm{\eta}}=\mathbb{E}[\partial\bm{\psi}(\widehat{f},{\bf X};\bm{\eta})/\partial\bm{\eta}]$, and $\rho=\dfrac{n}{N}$.
@@
$\bm{\omega}=\bm{\omega}_{0}$ $=(\operatorname{Var}(\widehat{\bm{\eta}}_{\cal L})+\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U}))^{-1}\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ (42)
@@
By the theory of PSPA, PPI++, and PPI paper [35, 4, 3], we have
$n^{1/2}(\widehat{{\bm{\theta}}}_{\texttt{PSPA}}-{\bm{\theta}}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},\bm{\Sigma}({\rm diag}(\bm{\omega}^{*}_{\text{ele}})))$ (52)
@@
Denote ${\bf C}={\bf A}_{{\bm{\theta}}^{*}}^{-1}{\bf A}_{\bm{\eta}}$, we have PSPA, PPI++, and PPI is asymptotically equivalent to ${\rm diag}(\bm{\omega}^{*}_{\text{ele}}){\bf C},{\rm diag}(\bm{\omega}^{*}_{\text{tr}}){\bf C}$, and ${\rm diag}({\bf 1}){\bf C}$, respectively. This completes the proof. ∎
@@ A.3 Proof of proposition 2
Appendix B An example for understanding the difference between PSPS and PPI++
Consider a linear regression with two predictors: $Y\sim\theta_{1}X_{1}+\theta_{2}X_{2}$. The summary statistics for PSPS can be expressed as: $\left[\hat{\theta}_{1{\cal L}},\hat{\theta}_{2{\cal L}},\hat{\eta}_{1{\cal L}},\hat{\eta}_{2{\cal L}},\hat{\eta}_{1{\cal U}},\hat{\eta}_{2{\cal U}}\right]^{T}$ from linear regression analysis in labeled and unlabeled data.
For PSPS, since $\hat{\theta}_{\textnormal{{PSPS}}}=\left[\begin{array}[]{l}\hat{\theta}_{1{\cal L}}\\ \hat{\theta}_{2{\cal L}}\end{array}\right]-\left[\begin{array}[]{cc}w_{1}&w_{12}\\ w_{12}&w_{2}\end{array}\right]\left[\begin{array}[]{l}\hat{\eta}_{1{\cal L}}\\ \hat{\eta}_{2{\cal L}}\end{array}\right]+\left[\begin{array}[]{cc}w_{1}&w_{12}\\ w_{12}&w_{2}\end{array}\right]\left[\begin{array}[]{l}\hat{\eta}_{1{\cal U}}\\ \hat{\eta}_{2{\cal U}}\end{array}\right]$, the final estimator for $\theta_{1}$ is $\hat{\theta}_{\textnormal{{PSPS}},1}=\hat{\theta}_{1{\cal L}}-w_{1}\hat{\eta}_{1{\cal L}}+w_{1}\hat{\eta}_{1{\cal U}}-w_{12}\hat{\eta}_{2{\cal L}}+\omega_{12}\hat{\eta}_{2{\cal U}}$.
In comparison, since $\hat{\theta}_{\texttt{PPI++}}=\left[\begin{array}[]{l}\hat{\theta}_{1{\cal L}}\\ \hat{\theta}_{2{\cal L}}\end{array}\right]-\left[\begin{array}[]{ll}w&0\\ 0&w\end{array}\right]\left[\begin{array}[]{l}\hat{\eta}_{1{\cal L}}\\ \hat{\eta}_{2L}\end{array}\right]+\left[\begin{array}[]{cc}w&0\\ 0&w\end{array}\right]\left[\begin{array}[]{l}\hat{\eta}_{1{\cal U}}\\ \hat{\eta}_{2U}\end{array}\right]$, its estimator for $\theta_{1}$ is $\hat{\theta}_{\texttt{PPI++},1}=\hat{\theta}_{1{\cal L}}-w\hat{\eta}_{1{\cal L}}+w\hat{\eta}_{1{\cal U}}$.
Since $\hat{\theta}_{\textnormal{{PSPS}},1}$ involves two zero-augmentation terms (i.e., $-w_{1}\hat{\eta}_{1{\cal L}}+w_{1}\hat{\eta}_{1U}$ and $\left.-\omega_{12}\hat{\eta}_{2{\cal L}}+\omega_{12}\hat{\eta}_{2{\cal U}}\right)$, its asymptotic variance should be less than or equal to that of PPI++ with one augmentation term. Therefore, PSPS borrows information from both coordinates, but PPI++ is restricted to information from only the first coordinate. Although the PPI++ can be used under a different scalarization, it still contains one augmentation term.
Appendix C Algorithms for ML-assisted FDR control
Algorithm 2 PSPS-BH for linear regression
Algorithm 3 PSPS-knockoff with debiased Lasso
Here, we employ second-order multivariate Gaussian knockoff variables for knockoff-sample and use the difference between the absolute values of the $k$-th feature and its knockoff coefficient as the knockoff-statistic. Alternative choices for these two steps can also be readily integrated into our algorithm [26].
Appendix D Implementation details
D.1 Simulation
@@
Here, we provide the details for our simulation. All our simulation is run in R with version 4.2.1 (2022-06-23) in a MacBook Air with an M1 chip. For all the simulations, the ground truth coefficients are obtained using $5\times 10^{4}$ samples; A pre-trained random forest with 500 trees to grow is obtained from hold-out data. We bootstrap the labeled data for $200$ times for covariance estimation. All simulations are repeated 1000 times.
@@
  • Mean estimation, Linear regression, and Quantile regression: We simulate the data from $Y_{i}=X_{1i}\beta_{1}+X_{2i}\beta_{2}+X_{1i}^{2}\beta_{2}+X_{2i}^{2}\beta_{2}+\epsilon_{i}$, where $X_{i1}$ and $X_{2i}$ are independent simulated from $\mathcal{N}(0,1)$, $\beta_{1}=\sqrt{0.08}$, $\beta_{2}$ is set to be the value such that $X_{2i}\beta_{2}+X_{1i}^{2}\beta_{2}+X_{2i}^{2}\beta_{2}$ explains $81\%$ of the outcome variance, and $\epsilon_{i}$ is simulated from a mean zero normal distribution with variance such that $\operatorname{Var}(Y_{i})=1$. We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with sample size $1000,2500,5000$, or $10000$.
@@
  • Logistic regression: We simulate the data from $Y_{i}=\mathds{1}(\tilde{Y}_{i}>\text{ median }(\tilde{Y}_{i}))$, where $\tilde{Y}_{i}=X_{1i}\beta_{1}+X_{2i}\beta_{2}+X_{1i}^{2}\beta_{2}+X_{2i}^{2}\beta_{2}+\epsilon_{i}$, where $X_{i1}$ and $X_{2i}$ are independent simulated from $\mathcal{N}(0,1)$, $\beta_{1}=\sqrt{0.08}$, $\beta_{2}$ is set to be the value such that $X_{2i}\beta_{2}+X_{1i}^{2}\beta_{2}+X_{2i}^{2}\beta_{2}$ explains $81\%$ of the outcome variance, and $\epsilon_{i}$ is simulated from a mean zero normal distribution with variance such that $\operatorname{Var}(\tilde{Y}_{i})=1$. We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with sample size $1000,2500,5000$, or $10000$.
@@
    We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with sample size $1000,2500,5000$, or $10000$. The $Z_{i}$ is used as a instrument for $X_{1i}$.
@@
    We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with sample size $1000,2500,5000$, or $10000$.
@@
    We use $X_{1i},\dots,X_{200i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 100 individuals, and unlabeled data with sample size $1500,2000,2500$, or $3000$.
@@
    where $\beta_{1}=\sqrt{0.01}$ to power simulation and $\beta_{1}=0$ for type-I error simulation, $\beta_{2}$ is set to be the value such that $\beta_{2}X_{2i}+\beta_{2}X_{2i}^{2}$ explains $81\%$ of the outcome variance, and $\tau_{\epsilon}$ is set to be the value such that $\operatorname{Var}(Y_{i})=0$.We use $X_{1i}$ and $X_{2i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with sample size $1000,2500,5000$, or $10000$.
@@
    The correlation $r$ is set to be 0.25. We then simulate the outcome $Y_{i}=\sum_{k=1}^{150}X_{ki}\beta_{k}+\epsilon_{i}$, where we randomly sample $15$ $\beta_{k}$ to be 0.15 and let all other remaining $\beta_{k}$ to be 0. $\epsilon_{i}$ is simulated from a mean-zero normal distribution with variance set to the value such that $\operatorname{Var}(Y_{i})=1$. We further generate $Z_{i}=0.9Y_{i}+\sum_{k=1}^{150}X_{ki}\theta_{k}+\gamma_{i}$, where $\theta_{k}=0.15$ for all $k=1,\dots,150$. We use $Z_{i}$ as features to predict the $Y_{i}$ in the random forest. We consider labeled data with 500 individuals, and unlabeled data with sample size $5000$.
  • knockoff: We used the same setting as described in the BH procedure above to generate the data. The only difference is that we set $\beta_{k}=0.5$ for features associated with the outcome and considered labeled data consisting of 100 individuals, along with unlabeled data comprising a sample size of 1000.
D.2 Identify vQTLs for bone mineral density
Our prediction pipeline comprises two components: prediction for unlabeled data and prediction for labeled data. To predict bone mineral density in unlabeled data, we first selected predictive features by 1) calculating the correlation of bone mineral density with 466 other variables (sample size > 200,000 from UK Biobank) using labeled data and 2) selecting the top 50 variables with the highest correlations as inputs for the SoftImpute algorithm [30] to predict bone mineral density in the unlabeled data. For the labeled data, we employ a similar approach but incorporate 10-fold cross-validation to prevent overfitting. We select the predictive variables and train the SoftImpute model using 90% of the labeled data. We then perform predictions on the remaining 10% in each fold and repeat this process 10 times across all folds.
Appendix E Supplementary figures and tables
Figure E.1: Simulation for FDR control. Panel a-b shows the estimated FDR level given the expected FDR. Panel c-d shows the power.
Figure E.2: Manhattan plot of vQTLs for bone mineral density. The X-axis represents chromosomes (CHR), plotted by base pair positions (BP). Each point on the plot indicates a single nucleotide polymorphism (SNP). The Y-axis depicts -log10(p-values).
CHR BP SNP A1 A2 EAF BETA SE p-value FDR
1 205359339 rs12139623 T C 0.105 0.063 0.012 7.2e-08 0.033
4 14359045 rs552582509 A G 0.213 0.046 0.009 4.5e-07 0.045
7 132568586 rs79089873 A G 0.073 -0.068 0.014 5.6e-07 0.049
9 11587905 rs146938822 A G 0.022 -0.134 0.024 4.0e-08 0.033
15 47672201 rs281258 C T 0.393 -0.04 0.007 5.8e-08 0.033
Table E.1: Significant vQTLs for bone mineral density. Abbreviations: CHR, Chromosome; BP, Base Pair; SNP, Single Nucleotide Polymorphism; A1, Allele 1 (Effect Allele); A2, Allele 2 (Non-effect Allele); EAF, Effect Allele Frequency; BETA, Effect Size (Beta Coefficient); SE, Standard Error; FDR, False Discovery Rate.
Method Linear regression Logistic regression
PSPS 1.62s 8.27s
PPI 0.024s 0.032s
PPI++ 0.031s 0.077s
PSPA 0.049s 0.034s
Table E.2: Runtime experiments. Utilizing a dataset with 500 labeled and 10,000 unlabeled data points, PSPS required 1.62 seconds for linear regression and 8.27 seconds for logistic regression using 200 bootstrap resampling. The computation of one-step debiasing using summary statistics alone took 0.032 seconds for linear regression and 0.033 seconds for logistic regression. Current methods, which estimate asymptotic variance via the closed form derived by the Central Limit Theorem instead of resampling, ranged from 0.024 to 0.049 seconds for linear regression and 0.032 to 0.077 seconds for logistic regression.